OpenStudy (anonymous):

A combustion device was used to determine the empirical formula of a compound containing only carbon, hydrogen, and oxygen. A 0.6349 g sample of the unknown produced 1.603g of CO2 and 0.2810g H20. Determine the empirical formula of the compound.

OpenStudy (abmon98):

Mass fraction: 12/44=0.27 0.27*1.603=0.437grams of C in CO2 Molar Mass of H2O:18grams Mass fraction of Hydrogen in H2O: 2/18=0.111 0.111*0.2810=0.03 grams in original compound: 0.6349-(0.437+0.111)=0.1678 grams of Oxygen 0.437/12=0.036 0.03/1=0.03 0.1678/16=0.01048 Divide by the smalles figure C3H3O